### v1

# Evaluating the Utility of Conformal Prediction Sets for AI-Advised Image Labeling

Abstract.
As deep neural networks are more commonly deployed in high-stakes domains, their lack of interpretability makes uncertainty quantification challenging. We investigate the effects of presenting conformal prediction sets—a method for generating valid confidence sets in distribution-free uncertainty quantification—to express uncertainty in AI-advised decision-making. Through a large pre-registered experiment, we compare the utility of conformal prediction sets to displays of Top-1 and Top-k predictions for AI-advised image labeling. We find that the utility of prediction sets for accuracy varies with the difficulty of the task: while they result in accuracy on par with or less than Top-1 and Top-k displays for easy images, prediction sets excel at assisting humans in labeling out-of-distribution (OOD) images especially when the set size is small. Our results empirically pinpoint the practical challenges of conformal prediction sets and provide implications on how to incorporate them for real-world decision-making.
1. Introduction
As machine learning models like deep Neural Networks (NNs) achieve impressive predictive performance, AI-advised decision-making has become more commonplace in various tasks where a human must make a decision, from high stakes domains like medicine (e.g., (Razzak et al., 2018)) or autonomous vehicle navigation (e.g., (Fujiyoshi et al., 2019)) to online tasks like content moderation (e.g., (Lai et al., 2022)) or image labeling (e.g., (van der Wal et al., 2021)). Ideally, having access to a predictive model can help humans to make more informed decisions, surpassing the capabilities of either human judgment or AI in isolation, a synergy known as complementary performance (Bansal et al., 2021; Hemmer et al., 2021). Whenever an AI system understands certain regions of the feature space better, the human counterpart can benefit by considering the model predictions. For example, in AI-advised image labeling, which we study in this paper, human labelers may benefit from viewing predictions from a model trained on previously labeled examples.
To help humans know when to trust model predictions, presenting information about the probability that the model’s prediction is correct can in principle be useful, such as providing the softmax pseudo-probability produced in the final layer of NNs for classification tasks. However, predictions from NNs are known to be overconfident (Guo et al., 2017), and heuristic measures of uncertainty are poorly calibrated. Consequently, decision-making relying on definitive NN predictions (e.g., Top-) with the highest softmax prediction confidence can lead to an unwarranted impression that prediction uncertainty is minimal or not important, affecting the quality of the decisions that are ultimately made.
Given these challenges, conformal prediction (Vovk et al., 2005; Romano et al., 2020; Angelopoulos et al., 2020) has emerged as an alternative solution that can rigorously quantify the prediction uncertainty for NNs through the use of prediction sets. Instead of presenting unreliable softmax values or other heuristics for uncertainty, the prediction set is a type of confidence interval, consisting of a set of predictions with a coverage guarantee—the true class is, on average, captured within such sets with a user-specified probability (e.g., ). The conformal framework is compatible with NNs because the derived sets are distribution-free—they are model-agnostic, work with finite samples, and offer non-asymptotic guarantees without distributional assumptions (Angelopoulos and Bates, 2021).
One challenge is that while conformal prediction sets express prediction uncertainty with guarantees that status quo presentations like Top- predictions with softmax values cannot, the conformal coverage guarantee can lead to large sets embodying high uncertainty for instances the model perceives as difficult (Babbar et al., 2022; Angelopoulos et al., 2020). Compared with Top- predictions, such sets may increase cognitive load, rendering uncertainty quantification less effective (Zhou et al., 2017), and thus compromising predictive performance. To understand how conformal prediction sets support human decisions aided by model predictions in an image labeling task, we contribute a large pre-registered online experiment that compares the accuracy achieved in AI-advised image labeling using conformal prediction sets with that achieved with no predictions or Top- predictions. Participants in our study were tasked with labeling images from ILSVRC 2012 (Deng et al., 2009) that varied in difficulty, and encompassed both image stimuli that were in-distribution and out-of-distribution (OOD). This allowed us to evaluate the utility of prediction sets against Top- presentation alternatives in scenarios where the model’s predictions are produced by a well-calibrated model and when predictions are unreliable (i.e., more error-prone), such as when models are poorly calibrated due to shifts in the distribution of new instances relative to training data.
We evaluate decision-making by metrics such as accuracy and the shortest path length in the label space hierarchy, which approximates the amount of deviation from the ground truth in the label space hierarchy. We also use participants’ elicited willingness-to-pay at the end of the experiment to compare the perceived value of predictions against the post-hoc inferred monetary benefits derived from accessing these predictions.
We find that for in-distribution instances, prediction sets lead to reduced labeling accuracy compared to Top- predictions. However, for OOD instances where the model misspecification differentially affected the calibration of the Top- predictions, prediction sets improve accuracy regardless of set size. We find that participants are willing to pay roughly equivalent amounts for each type of display; if anything, they undervalue prediction sets relative to other presentations despite their performance advantages. Our results advance understanding of the strengths and limitations of prediction sets as a more robust form of uncertainty quantification for improving AI-advised decisions, and highlight how decision-makers can over-rely on provided predictions, even when poorly calibrated, in an image labeling task.
2. Background
2.1. Uncertainty Quantification Using Conformal Prediction
Prior work on Human-Computer Interaction (HCI) suggests that humans can make more informed decisions when uncertainty is effectively communicated (e.g., (Joslyn and LeClerc, 2012; Kay et al., 2016; Hullman et al., 2018)), such as by presenting prediction uncertainty through static intervals (Cumming, 2014; Cumming and Finch, 2005; Manski, 2019; Taylor and Kuyatt, 1994) or animations (Hullman et al., 2015; Zhang et al., 2021). However, uncertainty quantification for NNs remains challenging (Amodei et al., 2016; Angelopoulos and Bates, 2021) due to their poor interpretability and often unreliable softmax pseudo-probabilities (Guo et al., 2017; Torralba and Efros, 2011).
One approach to quantify prediction uncertainty in classification tasks is to construct Bayesian NNs, so that predictions, which depend on sampling from the posterior distributions of model weights, carry an inherent, mathematically-grounded measure of uncertainty. However, Bayesian NNs heavily rely on distributional assumptions for their priors and can be computationally intensive to train. Conformal prediction (Vovk et al., 2005) has emerged as a more flexible and efficient solution that does not require a strong model or distributional assumptions. In classification settings, conformal prediction quantifies uncertainty through prediction sets. Similar to traditional confidence intervals, the derived set contains predicted classes and achieves a coverage guarantee that, on average, the true class is captured by the set with a user-specified probability. Formally, given a training set , where represents the feature vector and represents its corresponding label, the uncertainty set function, denoted , maps a feature vector to a subset of to ensure the coverage guarantee with an error rate of .
If an uncertainty quantification approach is well calibrated, we naturally expect it to reflect greater uncertainty on instances that are more difficult for the model (e.g., OOD instances). With prediction sets (Romano et al., 2020; Angelopoulos et al., 2020), this manifests as adaptiveness: conformal prediction set will be larger when the task is more difficult for the model. Adaptiveness is realized through the inductive conformal method, which quantifies the prediction uncertainty of a chosen classifier without changing the model using a designated calibration set (Papadopoulos et al., 2002), in which instances are independently and identically distributed (i.i.d.) as those on which the model is trained.
For each calibration instance, the predicted softmax scores are adjusted by Platt scaling (Platt, 1999; Guo et al., 2017; Nixon et al., 2019) for better interpretability and further regularized to penalize unlikely classes (Angelopoulos et al., 2020). These calibrated probabilities are then ranked in descending order, reflecting the classifier’s confidence in each class. A conformity score is computed by summing these ranked probabilities up to the true class, capturing the model’s accumulated confidence for the instance leading up to the correct classification. From the distribution of conformity scores of all calibration instances, a threshold is computed by taking the quantile, after a finite sample correction for robustness. This calibration threshold makes the prediction set adaptive and maintains the coverage guarantee. For a new and unknown instance, this calibration threshold decides the size of the prediction set by including all labels whose softmax scores exceed this calibration threshold. In our study, we opted for the Regularized Adaptive Prediction Sets (RAPS) algorithm (Angelopoulos et al., 2020) to ensure that the prediction sets of our experiment are as small as possible while achieving, on average, the desired coverage rate.
2.2. Effects of Presenting Uncertainty in AI-advised Decisions
Pairing humans with an AI system in complex decision-making scenarios can elevate performance beyond what either can achieve individually, especially when their strengths are complementary (Bansal et al., 2021; Steyvers et al., 2022). Particularly, in AI-advised image labeling, a plethora of tools and software exists for boosting human annotators’ speed and accuracy (Sager et al., 2021). However, performance can be influenced by how predictions are presented—specifically, whether uncertainty is effectively communicated or not. When uncertainty is well communicated, predictions appear to be more reliable and trustworthy to humans (Zhang et al., 2020; Lai and Tan, 2019; Beller et al., 2013). Knowing a model’s prediction accuracy or the data-generating process can make predictions more transparent and explainable, which encourages analytical thinking and reduces over-reliance on predictions (Zhang et al., 2020; Prabhudesai et al., 2023). However, uncertainty quantification does not guarantee improved decision-making. Cognitive biases, such as tendencies to under- or over-react to information relative to rational standards play a major role in how one interprets uncertainty (e.g., (Ambuehl and Li, 2018; Tversky and Kahneman, 1971)). Additionally, Zhou et al. (2017) suggest that uncertainty quantification can improve trust under low cognitive load but has been found to diminish trust when uncertainty is presented ambiguously, demanding a high cognitive load.
There is limited empirical evidence on how prediction sets influence trust and accuracy in AI-advised decision-making. One exception is Babbar et al. (2022), who investigated the perceived utility and performance of humans using RAPS versus Top- predictions for labeling images from CIFAR-100 (Krizhevsky, 2009). They found that participants using RAPS reported higher trust and perceived utility than those using Top- predictions. However, they found a negligible difference in performance between the two groups, which may be due to their small study size (i.e., 15 in each group) and a lack of variation in their image stimuli. They suggested that smaller sets might lead to better accuracy, but this observation was based on a comparison between sets generated by two different conformal methods. Recent work by Straitouri and Rodriguez (2023) compared labeling performance between two decision support systems: one that only allowed participants to select a label value from the prediction set and another that allowed them to choose freely one of 16 labels. Their results suggest that when the model is well calibrated, decision-making solely based on prediction sets tends to be more reliable and accurate because of the coverage guarantee.
In contrast to Babbar et al. (2022), we do a much larger experiment that evaluates the utility of different uncertainty prediction displays for in-distribution and OOD stimuli, each of which is systematically categorized by difficulty and set size of the associated conformal prediction set. This enables us to explore changes in performance by instance type. Instead of restricting agency as in Straitouri and Rodriguez (2023), we create a more realistic labeling scenario where participants can freely search for their preferred choice from a vast label space hierarchy, as inaccurate predictions may still be useful serving as clues for participants to identify the correct answer.
3. Online Experiment
3.1. Overview
We conducted a large pre-registered111Pre-registration: https://aspredicted.org/6gh27.pdf mixed-design repeated measure experiment on Prolific to evaluate how different displays of prediction uncertainty impact performance on an AI-advised labeling task, including relative to performance without access to predictions. Figure 1 provides an overview of our stimuli generation process, which systematically varied whether the image instances were in-distribution or OOD and easy or difficult, as well as varying the size of the prediction sets produced by conformal prediction.
Experiment diagram demonstrating the pipeline and manipulations.
Each participant in our experiment labels a set of 16 images, sampled from the ILSVRC 2012. We assign participants to one of four conditions representing different prediction displays: (1) no predictions (baseline), (2) Top- prediction with softmax, (3) Top- predictions with softmax, or (4) a prediction set (RAPS) (Angelopoulos et al., 2020).
In real-world scenarios, the deployed model can encounter unexpected OOD instances where the distribution of input features differs from what was trained and calibrated. We vary whether images are in-distribution or OOD within subjects. Among the 16 stimuli, 4 are in distribution, while 12 are not, as we expect greater variation in participants’ accuracy for OOD.
In our experiment, OOD instances notably lower the model’s Top- prediction accuracy. While in some scenarios conformal coverage guarantees can be compromised by OOD stimuli, for the corrupted images from the ILSVRC 2012 dataset that we use as OOD instances, the adaptiveness of RAPS leads to good coverage despite OOD. However, the size of the prediction sets in number of labels included becomes much larger when the instance is harder for the model. To evaluate the impact of set size on the effectiveness of prediction sets, we balance set size within subject by categorizing images as small or large based on the size of the prediction set generated for the image. This factor enables us to examine the performance of the prediction set between size levels, as previous work suggests that smaller set size is more preferable (Romano et al., 2020; Angelopoulos et al., 2020) and may yield higher accuracy for in-distribution images (Babbar et al., 2022). We expect small set sizes to be more useful in general, as they reduce cognitive load while maintaining the conformal coverage guarantee for in-distribution images. Large sets embodying high uncertainty may still be useful for labeling when participants are highly uncertain of what the image shows and consequently heavily reliant on prediction sets. In our setting, set size can also be viewed as a nuanced measure of difficulty for in-distribution stimuli because images that require a larger set size to achieve the desired coverage are intrinsically more difficult (Angelopoulos et al., 2020).
For both in-distribution and OOD stimuli, we balance difficulty within subject by easy and hard, using the median cross-entropy loss of in-distribution images as the threshold. The former enables us to assess whether participants are more inclined to accept predictions when they express high confidence. The latter allows us to examine whether participants are more likely to modify or reject low-confidence predictions and instead rely on their own intuition.
Our study evaluates participants’ performance by their labeling accuracy and the shortest path length between the chosen label and the correct label in the WordNet hierarchy (Miller, 1995; Fellbaum, 1998). After they labeled all images, we asked them to report their willingness-to-pay for the prediction display they used (a value ranging from to USD, the maximum they could earn) if they were to repeat the study with a similar distribution of instances. Finally, we elicited the strategies employed to approach tasks.
3.2. Task and Stimuli Generation
3.2.1. Overview of Stimuli Generation
To generate stimuli, we sourced 80 task images from ILSVRC 2012, equally distributed between in-distribution and OOD (i.e., 40 in each). Each set of stimuli is further categorized by difficulty (between easy and hard) and set size (between small and large). This results in four distinct groups, each containing ten images, for both in-distribution and OOD. Image stimuli in each group were selected to exemplify their respective categories, using the model’s (described below) prediction confidence and the size of the derived prediction set.
3.2.2. Predictive Model for Image Classification
Given that underperforming models are rarely deployed in real-world scenarios, we deployed a classifier designed to achieve high prediction accuracy. We use a pre-trained Wide Residual Network (WRN) (Zagoruyko and Komodakis, 2016) on ILSVRC 2012. The WRN is a Residual Network (ResNet) variant (He et al., 2016), and its primary architectural distinction is the widening factor in the channel width of convolutional layers compared to a ResNet of the same depth. Specifically, we utilize a wide_resnet101_2 model with layers and a widening factor of , using PyTorch’s IMAGENET1K_V2 weights (Paszke et al., 2019). This model is well calibrated for in-distribution instances—achieving a Top- accuracy of and a Top- accuracy of over all in-distribution images, as shown in Table S1 of Appendix A. We use WRN to create Top- and Top- predictions and their corresponding softmax pseudo-probabilities to establish prediction displays for Top- conditions.
3.2.3. Calibration Set
To produce prediction sets, we randomly sample half of the 50,000 images from the ILSVRC 2012 validation set to create a conformal calibration set, which contains 25,000 in-distribution images that are unseen by the model. This calibration set is used to perform inductive conformalization by calibrating the model based on the conformity scores derived from all images of the conformal calibration set, which enables the creation of prediction sets for the remaining 25,000 images in the conformal hold-out set. Our prediction sets achieve size-stratified conditional coverage (), as shown in Table 1. It implies that, on average, there is a probability that an image’s true label class falls within the prediction set, conditioned on all set sizes.
| Size | Count | Coverage |
| 0 to 1 | 13,967 | 0.961 |
| 2 to 15 | 6,849 | 0.946 |
| 16 to 30 | 1,583 | 0.978 |
| 31 to 45 | 874 | 0.975 |
| 46 to 60 | 548 | 0.974 |
| 101 to 1,000 | 1,179 | 0.974 |
Generating OOD images
We use OOD images to allow comparing the utility of different prediction displays in cases where the classifier is poorly calibrated, which lowers the accuracy of Top- predictions and increases the size of prediction sets. To create OOD images, we take a systematic approach to transform in-distribution images from our conformal hold-out set to OOD by applying covariate shifts (Quinonero-Candela et al., 2008) through synthetic image corruption (Michaelis et al., 2019). In supervised learning, where , the joint distribution over images and labels can be represented as , covariate shift refers to alterations to , the probability of the features , without altering the classifier .
After experimenting with 15 corruption types introduced by Michaelis et al. (2019), we opted to use five corruptions that reduced WRN’s Top- prediction accuracy the most. These corruptions are: (1) defocus blur, (2) snow, (3) frost, (4) zoom blur, and (5) glass blur, which are bolded in Table S1 of Appendix A. Given that the human visual system can be resilient to certain minor changes (Michaelis et al., 2019; Hendrycks and Dietterich, 2019), we corrupt all images in the conformal hold-out set at the maximum severity, as defined by Michaelis et al. (2019).
Table S1 shows that our prediction sets achieve a coverage rate for in-distribution images (Vovk et al., 2005; Shafer and Vovk, 2008). However, the coverage guarantee becomes invalid for OOD samples (Angelopoulos et al., 2020) as images used during calibration and testing are no longer exchangeable (i.i.d.).
Classifying stimuli by difficulty
To compare the effects of the different prediction displays on images of varying difficulty, we categorize images by difficulty using cross-entropy loss, which calculates how well the predicted probability distribution aligns with the ground truth distribution. Formally in Equation 1, let be an input image; denotes the true probability distribution of the image’s label in one-hot encoding, while is our classifier’s predicted distribution.
Figure S1 in Appendix A displays the distribution of cross-entropy loss of all in-distribution images, with the median presented as an orange dotted line. Due to WRN’s high predictive power, this distribution is highly right-skewed, which creates a problem: if we use this median threshold to categorize difficulty, we would include many images in the easy group that the model is both confident and accurate in predictions, so the derived prediction sets embody little uncertainty, synonymous to Top- prediction. We thus compute a new median threshold () after excluding images with a set size of one such that the difficulty threshold, shown as a red solid line, is defined based on images with non-deterministic sets. Images with cross-entropy loss falling below or on this threshold are classified as easy, while those above the threshold are deemed hard.
Because the cross-entropy loss derived from the predictive distribution of OOD images can be misleading as a signal of task difficulty due to model misspecification (Guo et al., 2017; Torralba and Efros, 2011), we use the red median threshold from the in-distribution data to categorize the difficulty of OOD images.
| Hold-out Set | w/ Size = 1 | w/o Size = 1 | ||
| Easy | Hard | Easy | Hard | |
| None | 1 | 15 | 5 | 19 |
| defocus blur | 2 | 59 | 8 | 60 |
| snow | 1 | 41 | 7 | 43 |
| frost | 2 | 46 | 8 | 50 |
| zoom blur | 2 | 44 | 10 | 47 |
| glass blur | 2 | 63 | 9 | 65 |
Classifying stimuli by set size
Images within each difficulty group of the conformal hold-out set can induce different prediction set sizes. For example, although images classified as easy are more likely to have a small set size, there are easy instances that require a large set size to achieve the desired coverage. Because we wish to understand how set size can impact the efficacy of prediction sets, we categorize stimuli within each difficulty group by size. After grouping images in each conformal hold-out set by easy and hard, we distinguish between images based on relative set size as produced by RAPS, categorizing them as either small or large, based on the median of the group. However, as shown in Table 2, the predictive power of WRN consistently demonstrates high confidence in its predictions for many easy images with a set size of one, which significantly lowers the median size threshold. To ensure more variance in set size within the easy groups, we compute the median threshold after excluding images with a set size of one (i.e., these are directly assigned to the small group). Images with a set size smaller than or equal to the size median are classified as small; if the opposite is true, they are classified as large. The specific thresholds used for each difficulty category within the conformal hold-out set are outlined in Table 2, with the bold figures representing the applied thresholds.
Sampling task images based on defined categories
Upon categorizing all images in each conformal hold-out set (i.e., one in-distribution and five OOD, totaling 6 25,000 images) by difficulty (easy and hard) and size (small and large), our objective is to select a set of task images from which to sample image assignments for participants, so as to avoid results that are overfit to a very small set of stimuli. Specifically, we target 80 task images from the 1,200 candidate images. Our goals are to ensure: (1) the candidate images are representative of each group defined by in-distribution or OOD, difficulty, and size (), and (2) task images avoid the known data anomalies in ILSVRC 2012 (Ekambaram et al., 2017; Northcutt et al., 2021).
Example stimuli used for the experiment.
Sampling images from groups, such as in-distribution, easy (difficulty), and small (set size), is challenging due to the prevalence of images with a set size of one, for which prediction sets become nearly identical to the Top- prediction. To avoid oversampling sets of size one, (1) we temporarily set aside images with a set size of one, (2) we take images with sets larger than one whose set size falls near the median size in that group (i.e., and percentiles), and (3) perform a post hoc sampling correction to include images with a set size of one in proportion to their original presence in each group, as shown in Table S2 of Appendix A.
Images from ILSVRC 2012 are known to be noisy, with numerous data anomalies. To create the set of 80 task images, we thereby sample 50 candidate task images from each representative subset defined by in-distribution or OOD, difficulty, and size () without replacement. We manually inspect the task images, starting from images with higher cross-entropy loss (hard), and omit those with (1) an obviously wrong ground truth label, (2) a text overlay with the correct label, (3) no obvious focal object or many surrounding objects, (4) unusual dimensions, and (5) small multiples. Based on these rules, we selected 10 images from each of the 4 groups generated by the combination of that are in-distribution, and 2 images from each of the 4 groups with corruptions. This process generates 80 task images in total, balanced by in-distribution or OOD, difficulty, and set size.
Table 3 presents the Top- accuracy, coverage, and mean set size grouped by our experimental factors with example stimuli shown in Figure 2. Our stimuli achieve the desired behavior, as the average prediction accuracy and set size vary according to the difficulty and set size manipulated. Notably, despite Top- prediction accuracy being lower for OOD stimuli, prediction sets for OOD stimuli maintain a high coverage rate. This is because the adaptiveness of the conformal prediction set (Romano et al., 2020; Angelopoulos et al., 2020) makes them more resilient to random perturbations from image corruptions. The prediction set can safeguard the coverage for certain instances by significantly increasing the set size, as shown by the bolded size summary in Table 3.
| Shift | Difficulty | Size | Count | acc@1 | acc@10 | Coverage | Avg. Set Size |
| in | easy | small | 10 | 1 | 1 | 1 | 2.6 |
| large | 10 | 1 | 1 | 1 | 19.5 | ||
| hard | small | 10 | 0.4 | 1 | 0.9 | 8.0 | |
| large | 10 | 0.5 | 0.9 | 0.9 | 51.1 | ||
| out | easy | small | 10 | 1 | 1 | 1 | 5 |
| large | 10 | 1 | 1 | 1 | 28.1 | ||
| hard | small | 10 | 0 | 0.6 | 0.9 | 30.1 | |
| large | 10 | 0.1 | 0.2 | 0.9 | 90.8 |
3.2.4. Reward and Bonus
Participants earn a base compensation of for completing the study and get a bonus for each correct label. With 16 trials, the maximum bonus is , bringing the total potential earnings to . We determined the task reward to achieve local minimum wage based on the assumption that most participants would take longer than the average time we observed in a small pilot of the experiment.
3.3. Experimental Interface
Our task interface222Study interface: https://conformal-prediction.com, of which screenshots are presented in Figure 3, uses a two-column design, presenting task images on the left and predictions on the right, with a response field directly below the task image. In the baseline condition without predictions, the task image and response field are centered. Participants can see real-time earnings in the top-right corner of the interface.
To ensure that participants understood the meaning of each label, we selected label-representative images from the conformal calibration set. To find suitable label-representatives, we manually examined all images by labels and prioritized selection for those with small cross-entropy loss (i.e., easy), following the same rules used when selecting task images. The label-representative images we used are included in the Supplemental Material.
3.3.1. Presenting Predictions
For the three treatment conditions that feature predictions, the Top- condition presents participants with the classifier’s Top-1 prediction and its softmax score on top of a representative image of the predicted label. Top- predictions with softmax scores are ordered in a table, while the prediction set is presented in an -by- matrix sorted column-wise by the classifier’s prediction confidence. Participants could hover their mouse cursor over any predicted label to view a label-representative image.
Interface screenshots by conditions.
3.3.2. Label Space Search
To label task images, participants need to be aware of the label space. Because ILSVRC 2012 has a large label space of 1,000, which is cumbersome to present in a dropdown or checkbox format, we organized labels using a version of the WordNet hierarchy, which we truncated slightly to avoid label overlap. For example, automobile and truck are both parent nodes for minivan, and since truck is also a type of automobile, we keep minivan in the automobile category. We include this tree in the Supplemental Material.
If participants disagree with the predictions or when no prediction is available (baseline), our interface enables them to perform three types of search. As illustrated in Figure 4, they are: (A) dropdown search, (B) keyword search, and (C) bottom-up search. In dropdown search, a dropdown menu appears as participants begin typing in the response field, suggesting leaf labels that match their input. Alternatively, in keyword search, participants can type and then click a search button to search the hierarchy for specific keywords. Doing so opens an overlay window showing the label-space hierarchy network, where leaf nodes (denoting valid labels) appear in a different color from the parent nodes. Clicking on a leaf node populates the label in the response field. To ease navigation, the hierarchy network is collapsed by default except for the parent nodes (i.e., categories) and leaf nodes (i.e., labels) containing the search keywords with edges connecting them highlighted in orange. All other parent nodes can also be clicked to expand.
Participants can also perform a bottom-up search starting from the prediction display. Clicking on a predicted label opens the search overlay. Instead of showing the whole hierarchy network, the overlay presents the path from the root node to the leaf node corresponding to the predicted label the participant clicked. Participants can explore adjacent labels under the same parent category or explore other parent nodes on the path upward toward the root node.
When exploring the label space hierarchy, participants can hover over any leaf node to display its label-representative image. We illustrate these search actions through video demonstrations included in the Supplemental Material.
Interface screenshots by conditions.
3.4. Experimental Procedure
Participants are directed to our study interface from Prolific. On the welcome page, they receive a brief description of the study’s purpose and the estimated time required for completion. Participants are instructed to complete the study in a single session using Google Chrome on a large-screen device. If participants agree to the terms by clicking the “Start” button, they are randomly assigned to a prediction display condition with task images sampled from each corresponding group, shuffled in a randomized order.
Participants then review two instruction pages. The first introduces the interface layout and details the bonus mechanism. The second contains several short videos, and demonstrates how to explore predictions (except in the baseline condition) and navigate the label space. Participants must watch all provided instructional videos before proceeding to an example page where they can practice using the interface. Then, they complete 16 rounds of tasks. Upon finishing the 16 trials, participants are asked to state their willingness-to-pay to have access to the same style of prediction display they used in the experiment on a scale from to USD (i.e., the range of possible earnings from the experiment) if they were invited to undertake another 16 rounds of tasks with new images generated similarly to those they had experienced, for further reward and potential bonus. On the same page, participants could optionally describe their strategy to identify the correct label for each image, with or without aid from predictions.
3.5. Participants
We recruited 600 participants from Prolific, divided equally among the four prediction displays, ensuring a gender-balanced sample. We set additional screening criteria so that all of our participants are (1) based in the USA; (2) no vision impairment; (3) no colorblindness; (4) English as the first language; and (5) aged between 18 and 65. This led to 150 participants in each condition representing different prediction displays. We determined the target sample size using a nonparametric simulation of the experiment bootstrapped from a pilot study of 30 participants. By grouping trial-level responses according to all manipulated factors and bootstrapping these responses with replacement, we simulated trial-level outcomes. We identified the target sample size (600) that resulted in confidence intervals on accuracy with width less than .
3.6. Analysis Method
We pre-registered an analysis plan focusing on three response variables: (1) accuracy, (2) shortest path length, and (3) willingness-to-pay, and followed our pre-registered plan in all analyses reported below unless otherwise specified. We define accuracy (ACC) as binary, indicating whether a participant selected the correct label for the task. Shortest path (SP) length is a continuous variable representing the number of hops from the leaf node a participant selects to the leaf node of the ground truth label. SP length is a measure widely used to quantify the semantic similarity (Resnik, 1995; Budanitsky and Hirst, 2006) between categories in a taxonomy. Although highly correlated with accuracy, this measure provides further information on how “incorrectness” varied by our experimental manipulations. Willingness-to-pay (WTP) ranges from and , and measures how much they will pay to have access to the display.
We pre-registered Bayesian Linear Mixed-effect models for accuracy and shortest path length. Our model specifications are presented in Model 1 and Model 2. In this context, trial is numeric, indicating the trial number, whereas condition (4 levels: baseline, Top-, Top-, and prediction set), shift (2 levels: in-distribution and OOD), difficulty (2 levels: easy and hard), and size (2 levels: small and large) are factors corresponding to our experimental manipulations. The term ID represents each participant’s unique Prolific ID.
For the accuracy model, we use a Bernoulli distribution for the likelihood shown in Line 1 left, parameterized by , which denotes the probability of a correct label. Line 3 presents the hierarchical logistic regression model. In this specification, we model the logit of through interactions among all fixed effects and incorporate random intercepts and slopes for trial, grouped by participants’ unique ID. This approach allows us to account for participants’ baseline performance differences and variations in rates of change across conditions and trials.
For the shortest path length model, we employ a Zero-inflated Negative Binomial (ZINB) model as the likelihood, parameterized by (expected SP length), (negative binomial dispersion), and (zero-inflation probability), as shown in Line 1 right. The expected path length between the participants’ responses and the truth label is modeled by the negative binomial component in Line 3. We apply a log-link function to , modeled by all fixed effects along with their interactions with random intercepts and slopes for trials grouped by participant ID. We do not explicitly model the dispersion parameter for the NB component. However, we placed an exponential prior with a rate of to moderate the degree of dispersion in our model. Given that our observed data had a maximum SP length of 27, this prior acts as a regularizing component, ensuring our model captures the dispersion of the observed data without predicting implausibly large SP lengths. Line 5 shows the zero-inflation component that models the proportion of zero path length (i.e., correct answer) not captured by the NB components. We apply a logit link function to and model it by all fixed effects along with their interactions.
We deviated from our pre-registered model specifications only in setting the priors. We originally aimed to set weakly informative priors, but several were too narrow to allow model convergence. We therefore widened the priors to be even less informative. Full details are available in supplemental material.
We followed a standard Bayesian workflow (Gelman et al., 2020) to check if both models fit. Detailed model diagnostics are included in the Supplemental Material. We use the models to generate predictions of the respective response variable based on 1,000 draws from the posterior distribution of the models’ fixed effects for each unique combination of our experimental manipulations while holding the trial effect constant at the mean. Our results present the expected predictions as point estimates, with uncertainty expressed through highest posterior density interval (HDI).
We pre-registered an analysis plan to assess participants’ reported WTP, relative to the aggregate observed value of each prediction display. Unlike traditional Likert-style questions, WTP allows us to define an objective benchmark post-hoc for comparison to participants’ responses.
Specifically, we first calculate the expected bonus participants gained in each condition by multiplying the expected accuracy rates by the maximum earnable bonus of . We then calculate the expected bonus difference with access to each prediction display relative to the baseline by subtracting the expected bonus of the baseline from the expected bonus of each treatment conditions. We finally evaluate the discrepancy between this expected bonus difference and the corresponding WTPs reported by participants who used each prediction display, by constructing a “willingness-to-overpay” metric. This measure reflects the difference between the additional bonus one is expected to earn from having access to a prediction display over the baseline and the actual difference in the average bonus participants received from accessing a prediction display relative to the baseline. We use bootstrapping to derive 95% confidence intervals with the median as a point estimate.
3.7. Qualitative Analysis of Strategies
We performed a qualitative analysis and developed a bottom-up open coding scheme using the strategies provided by the participants. Two coders worked in collaboration, with each coding half of the strategies. After coding, any ambiguous strategies were discussed until a mutual agreement was reached. Among 585 valid strategies, we excluded 41 uninformative ones (Code: 1) with responses such as “I have no idea, trying my hardest”. For the remaining strategies, we identified 17 ways in which the prediction displays and the search features were used. They are summarized by the categories below.
How do participants use different prediction displays? (Code: 2–6)
Do participants find predictions to be trustworthy? (Code: 7–10)
How do participants use the search feature to find a preferred label? (Code: 11–15)
How do participants narrow down their choices and identify a preferred label? (Code: 16–18)
4. Results
4.1. Data Preliminaries
We received 599 valid responses after excluding one participant from the baseline condition according to our pre-registered exclusion rule. As demonstrated in Figure 5 left, on average, participants spent minutes completing all 16 image labeling tasks for all conditions. Participants in Top-10 and RAPS conditions took slightly longer to complete than those of the baseline and Top-1 conditions.
Distribution of participants’ study completion time and the amount of bonus earned.
From the distribution of earned bonus shown in Figure 5 right, we see a clear difference that having access to prediction displays can help participants get more correct answers and hence a higher bonus reward than those who did not have access to the predictions. On average, baseline participants provided seven correct labels, which is lower than those who used prediction displays (10 out of 16 correct).
4.2. Accuracy
4.2.1. Overview
In Figure 6, we present the expected predictions by our accuracy model (Model 1) for each type of image stimuli, with uncertainty quantified as HDIs. At a high level, we find that RAPS does not improve participants’ labeling accuracy for easy tasks, but is more useful for hard tasks, especially for OOD images. While a smaller RAPS can be more useful for in-distribution images, size does not appear to affect accuracy much when images are OOD. Because small and large set sizes depend on whether instances are in-distribution or OOD in our data-generating model (Section 3.2), we provide the mean set size for size groups to contextualize results below.
Accuracy results
4.2.2. In-distribution
For both easy and hard images that are in-distribution, a larger set size may decrease participants’ labeling accuracy
This effect is most prominent when images are classified as hard in Figure 6, (C) and (D). When using RAPS to label hard images, participants’ accuracy is higher when set size is smaller (; HDI: ; mean size: ), and lower when set size is larger (; HDI: ; mean size: ). We speculate the reason is that a larger set size increases participants’ cognitive load as they need to traverse through the set and evaluate all labels. Figure 6 (C) shows that expected accuracy of RAPS participants is roughly % higher than that of the Top-10 condition (; HDI: ) which is roughly % higher than the Top-1 condition (; HDI: ), though HDIs overlap.
When the true label is in the prediction display, participants do not always choose it.
This is particularly evident for in-distribution images categorized as easy (cf. Figure 6, (A) and (B)). Here, the participants’ labeling accuracy tends to be lower than the prediction display accuracy (i.e., the percentage of the time that the prediction display included the true label), depicted as solid horizontal lines. However, when labeling hard images, participants in the Top-1 condition can improve upon the model’s prediction. As suggested by our open codes discussed in Section 4.5, we find when the Top-1 prediction was obviously inaccurate, participants were more likely to rely on their own judgment and used the search tool to identify the correct answer. The presence of a greater number of labels in the Top-10 and RAPS conditions may have complicated the labeling process due to increased cognitive load. This is exemplified by the observed trends in these conditions: among incorrect responses, participants in the RAPS condition missed the correct label from the prediction set of the time, whereas those in the Top-10 condition did so of the time.
In summary, when task images are in-distribution, (1) labeling accuracy for participants in the Top-1 or Top-10 conditions was higher when the images are easy (cf. Figure 6 (A) and (B)); (2) the accuracy of RAPS participants is negatively correlated with set size: a larger set size decreases accuracy, particularly for hard images (cf. Figure 6 (C) and (D)); and (3) Top-1 participants are more likely to rely on their own judgment and use the search tool to identify the correct answer when predictions were less likely to be correct. In contrast, the pattern of incorrect responses observed in the Top-10 and RAPS conditions suggests that the presence of more predictions in the display made it harder for participants to discern between inaccurate and accurate predictions, even when the correct label was present.
4.2.3. Out-of-distribution
When OOD images are easy, Top- predictions yield higher accuracy than RAPS
Due to the similarity in HDIs across conditions for easy OOD images (cf. Figure 6 (E) and (F)), we report expected accuracy marginalized over set sizes. We find Top-1 participants tend to achieve slightly higher accuracy in expectation (; HPI: ) than those using Top-10 (; HDP: ). Meanwhile, the RAPS condition yielded the lowest expected accuracy of (HPI: ), despite an average set size of (as detailed in Table 3) that is smaller than the size of Top-10, with identical chance that the prediction display contained the true label (colored lines in Figure 6 (E) and (F)). This divergence from our in-distribution finding, where smaller set sizes typically correlate with higher labeling accuracy when predictions are reliable, suggests additional factors may be at play in the OOD scenario. It is possible the reduction in cognitive load afforded by access to softmax pseudo-probabilities in the Top-10 condition for use in prioritizing labels aided participants in finding the correct label. This observation partially reinforces our finding from the easy in-distribution results: when the model is well-calibrated (cf. Figure 6 (E) and (F)), viewing fewer predictions and having access to prediction confidence (i.e., as in our Top- conditions) can improve accuracy.
When OOD images are hard, RAPS yield the highest labeling accuracy regardless of set size
For hard OOD images, RAPS set sizes grow substantially larger, with small sets averaging instances and large sets averaging instances. When hard OOD images had smaller set size (Figure 6 (G)), RAPS participants outperformed those in Top- conditions with a relatively high accuracy of (HPI: ), followed by the Top-10 condition (; HPI: ) and Top-1 condition (; HPI: ). When set size was larger (Figure 6 (H)), we identified a consistent pattern that RAPS participants (, HPI: ) consistently outperformed those in the Top- conditions. Nevertheless, the accuracy of RAPS participants is significantly lower than the average coverage rate (Figure 6 (G) and (F) green solid horizontal lines), implying that even when the true label is included in the set, it is still challenging for participants to identify it.
When OOD images are hard, participants relying on their own judgment may achieve better accuracy than those relying on poorly-calibrated Top- predictions
For hard OOD images with a smaller average set size (Figure 6, (G)), we find that participants in the baseline condition who did not have access to any predictions have a higher accuracy of (HPI: ) than those in the Top-1 condition (; HPI: ). For hard OOD images with a larger average set size (Figure 6, (H)), the accuracy of using RAPS (, HPI: ) versus the baseline condition (; HPI: ) is similar, and participants who used Top-1 prediction are more accurate (; HPI: ) than those who used Top-10 predictions (; HPI: ). Our results suggest that Top- predictions from a poorly calibrated model (cf. Figure 6 (G) and (H)) can negatively influence human judgment. Although RAPS participants had relatively higher accuracy than those in the Top- conditions, their accuracy is almost identical to those in baseline, suggesting it might be beneficial to withhold predictions for hard OOD stimuli that require a large set size to achieve the desired coverage.
In summary, (1) when OOD images are easy and the predictions and softmax scores well-calibrated (cf. Figure 6 (E) and (F)), Top- predictions tend to result in higher labeling accuracy; (2) when OOD images are hard and the model predictions more error-prone (cf. Figure 6 (G) and (H)), RAPS participants can outperform their Top- counterparts. Unlike in the in-distribution case, increasing the size of the prediction set does not degrade accuracy; and (3) when OOD task images are hard and the predictions more error-prone, participants who rely on their own judgment and the label space search tool may be more likely to identify the correct answer. The reasons behind the performance advantage of participants in the baseline condition when labeling hard OOD images with larger set sizes remain ambiguous. It is possible the accuracy decrease for participants in Top-10 condition when labeling hard OOD images that have larger set size is due to overreliance on inaccurate predictions, given that the average coverage for Top-10 predictions decreases from (small) to (large), as shown in Table 3.
4.3. Shortest Path Length
The predictions of our SP length model (Model 2) highly correlate with those of the accuracy model when comparing the results in Figure 6 with those in Figure 7. We briefly highlight the main takeaways that complement our accuracy results, with more detailed descriptions in Appendix B.1.
Shortest path length results
When OOD images are hard, a larger prediction set leads to submitted labels that are slightly closer to the correct label
Recall that in Section 4.2.3, we find when OOD images are classified as hard, RAPS participants can outperform their Top- counterparts, but there is no difference in accuracy by set size (cf. Figure 6, (G) and (H)). However, from the SP length, we find if the task image has a larger set size, RAPS participants’ chosen label is slightly closer to the correct answer in the label space hierarchy. As demonstrated in Figure 7 (G) and (H), the SP length for hard OOD images that have a larger set is (HPI: , mean size: ) and that of a smaller set is 4.2 (HPI: ; mean size: ). It is difficult to say why this might be the case; it is possible, as supported by our open codes in Section 4.5, that when OOD stimuli are especially challenging to recognize, participants spend more time scrutinizing labels in the prediction set clues, and that even when the probability that the prediction set contains the true label is the same, viewing more predictions helps cue their ability to identify the true label. The effect is quite small, whatever the cause.
When predictions are generated by a well-calibrated model, having access to predictions can guide participants closer to the correct answer
This result is most prominent for hard in-distribution images with smaller set size. From Figure 6 (C), we see the baseline and Top-1 prediction tend to yield similar accuracy, as HDIs closely overlap. However, in Figure 7 (C), we find that participants who used Top-1 prediction can provide responses that are closer to the correct answer than those in the baseline.
In summary, we find (1) for hard OOD stimuli, although the labeling accuracy of RAPS participants is similar, a larger prediction set may guide responses closer to the ground truth (cf. Figure 7 (G) and (H)); and (2) when the model is well calibrated, inaccurate Top- predictions might still guide participants to identify the correct label.
4.4. Willingness-to-Pay
We present the distributions of elicited willingness-to-pay (WTP) measures in Figure S2 with summary statistics in Table S4 of Appendix B.2. Elicited WTPs across conditions are quite similar, suggesting the measure may be noisy. Participants are willing to pay slightly more for Top-10 (; CI: ) than RAPS (; CI: ) and Top-1 (; CI: ).
Nonetheless, as detailed in Section 3.6, we quantify participants’ perceived utility for each prediction display using a measure we devised, called “willingness-to-overpay”. This proxy focuses on economic valuation, offering direct monetary comparisons between the monetary value of a prediction display that participants perceive (i.e., willingness-to-pay) against how much additional expected bonus participants gained from having access to each prediction display.
| Condition | Accuracy | Expected Bonus | Expected Bonus Difference |
| baseline | 0.41 | 1.64 | 0.00 |
| Top-1 | 0.63 | 2.52 | 0.86 |
| Top-10 | 0.64 | 2.56 | 0.88 |
| RAPS | 0.66 | 2.64 | 0.98 |
Bootstrapped 95% confidence interval for willingness-to-overpay
We note that this measure is imperfect as a comparator for any individual willingness-to-pay, because we cannot observe the counterfactual performance without a prediction display for each participant individually. In aggregate, however, Figure 8 suggests that participants may over-value the prediction displays to a similar extent across conditions. If anything, we see weak evidence that RAPS participants, despite benefiting more from using the prediction sets, perceive a lower monetary value from the prediction set, relative to Top- predictions. RAPS participants are willing to overpay by on average (CI ), slightly less than Top-1 participants (, CI ). Participants find Top-10 prediction most useful on average (; CI ).
4.5. Qualitative Analysis of Strategies
From our open coding, we find that participants employed different strategies between prediction displays (see Figure S3 in Appendix B.3). In general, most participants reported consulting predictions when decision-making ( out of 450). Some participants reported first inspecting predictions and assessing their accuracy using their judgment (Top-1: out of 150; Top-10: out of 150; RAPS: out of 150), while others report relying first on their own intuitions to form a general idea, and then seeing if it aligned with any predictions (Top-1: out of 150; Top-10: out of 150; RAPS: out of 150). Very few participants chose to rely completely on their own intuitions because they find predictions inaccurate or not useful (Top-1: out of 150; Top-10: out of 150; RAPS: out of 150). Others reported they would only consult predictions when the image was too hard to recognize and they had no clue (Top-1: out of 150; Top-10: out of 150; RAPS: out of 150). Among participants who reported that they trusted AI prediction(s), some of them in Top- conditions placed higher trust on predictions with high softmax pseudo-probability (e.g., ) (Top-1: out of 23; Top-10: out of 35), while those in RAPS condition focused more on the top suggestions in the prediction set ( out of 26).
Collectively participants used all features of our interface, reporting that the dropdown ( out of 599), bottom-up ( out of 599), and keyword search ( out of 599) methods were useful in helping them identify a label, and the representative images were informative for aiding understanding of the meaning of labels ( out of 599). Of participants who viewed prediction displays, those in the Top-1 condition appeared most likely to use the label space hierarchy network to find a better match ( out of 150, compared to out of 150 and out of 150).
In addition to our open codes, we summarized the proportion of participants who fully relied on predictions (i.e., with all submitted responses selected from predictions). Aligning with the codes described above, we find Top-1 participants relied less on the predictions and used the search tool to find their preferred choice more often for both in-distribution ( out of 150) and OOD ( out of 150) images. Top-10 participants reported relying heavily on predictions for in-distribution images ( out of 150) and much less for OOD ( out of 150).RAPS participants similarly reported very frequently basing their decisions on a constrained set of predictions in-distribution ( out of 150) and slightly less but still quite often for OOD ( out of 150).
5. Discussion
Conformal prediction has achieved recent visibility as a distribution-free approach to quantifying uncertainty in model predictions, which is robust to some forms of model misspecification. Our results suggest that decision-makers can use prediction sets effectively in an AI-advised labeling task, but that their benefits over status quo Top- presentation displays depend on the properties of the setting. In real-world scenarios where the deployed model is typically well calibrated and the predictions are reliable, we find the size of the prediction set is a critical determinant of its utility. Uncertainty quantification by prediction sets appears to be more useful with a smaller set size and less effective with a larger one that can increase users’ cognitive load. In contrast, when the model encounters unforeseen and challenging OOD instances that can cause Top- predictions and their associated uncertainty scores unreliable, prediction sets have a comparative advantage, regardless of their set size, at least in a setting like ours where coverage guarantees are maintained. We elucidate our main findings and discuss the broader implications of our study.
A smaller prediction set derived from a well-calibrated model is generally more useful.
Our results suggest that the utility of conformal prediction sets is linked to set size, presumably due to an underlying factor of cognitive load. A large set embodying more uncertainty requires users to navigate and evaluate many incorrect predictions before finding the correct answer. This increase in cognitive load makes prediction sets less effective and may encourage more satisficing (Simon, 1956), resulting in low response accuracy. Our results show that labeling accuracy using prediction sets is higher than those of Top- only when the uncertainty of in-distribution stimuli can be quantified by a smaller set. Despite earning higher bonuses, there is some evidence that participants who used prediction sets perceived lower worth compared to their Top- counterparts, as captured by their willingness-to-pay for the prediction displays. Therefore, uncertainty quantification does not guarantee more informed decision-making from users. We think designers of prediction displays should carefully consider the trade-off between set size and adaptiveness when using conformal prediction sets as a method to communicate prediction uncertainty for machine learning models in practical settings.
Conformal prediction sets are more useful for hard OOD instances.
One possible reason why prediction sets can exceed in performance for hard OOD stimuli is their adaptiveness. While the covariate shifts used in our experimental setting can severely affect the accuracy of Top- predictions, adaptive prediction sets under certain shifts can retain coverage by significantly increasing the set size (e.g., (Gendler et al., 2022)). While, the high coverage of prediction sets for OOD instances does not necessarily predict high accuracy of the AI-advised human, as the results for hard instances in Figure 6 demonstrate, even large set sizes for OOD instances provided some value to participants. Even when participants did not select the correct label from the set, when given a larger prediction set with hard OOD stimuli, participants tended to select labels that were closer to the correct one as measured by the shortest path length in the label space hierarchy. Hence, our findings highlight the value of conformal prediction sets in enhancing the reliability of predictions by machine learning models, especially when encountering OOD instances in real-world applications. Prediction display designers should leverage these sets’ adaptiveness and distribution-free properties.
Withholding predictions of an uncalibrated model can improve decision quality
Consistent with prior work on AI-advised decision-making (e.g., (Green and Chen, 2019; Beede et al., 2020; Carton et al., 2020; Liu et al., 2021)), our results suggest that when a model is well-calibrated and more accurate than humans making judgments alone, users with access to its predictions can perform better than they would without the model, but not as well as the model alone for easier instances. When the model is poorly calibrated, we find that what sort of prediction display is shown affects whether people can do better by accessing model predictions. For instance, when given access to Top- displays with low coverage of the true label, users sometimes did worse than they would have by ignoring the display. We observed qualitative feedback such as “I used the Top- when it was impossible to find what the image was based on my own perception” and “in cases where I really was unsure and did not know what the picture was, I relied on the AI and went with their label”. The quality of decision-making when users are presented with a constrained set of label options is highly correlated with the magnitude of model calibration and hence performance can vary with the calibration of the uncertainty quantification. Only those who used RAPS reliably met or exceeded the expected performance without any prediction display. Our results underscore the pivotal role calibration plays and the potential value of treating the detection of instance type as a model’s learning problem (e.g., (Kaur et al., 2023)).
5.1. Limitations and Future Work
It is unclear whether a larger prediction set showing more possible labels can still be effective if the coverage is reduced. A natural follow-up study is to evaluate the efficacy of prediction sets in an OOD setting where coverage is disrupted at varying magnitudes. The adaptiveness of prediction sets (Romano et al., 2020; Angelopoulos et al., 2020) enhances a model’s resilience to random and common perturbations that can occur in real-world scenarios, such as image corruptions. However, future work may explore the use of intentional adversarial methods that can attack softmax (e.g., FGSM (Goodfellow et al., 2015) or PGD (Madry et al., 2017)). Combined with the finding from Straitouri and Rodriguez (2023) that constrained choices can lead to better performance based on sets derived from a calibrated model with a smaller size, we might see an opposite effect than what we observe, where people are less likely to pick a choice from a prediction set that has a lower chance of containing the true label.
We think willingness-to-pay is an imperfect comparator because our experimental design does not observe the performance of individual participants without a prediction display. Future work can further explore how users’ perceptions of the value of prediction displays and compare performance and perceived value within-subject. Additionally, it is unclear whether the coverage guarantee of conformal prediction sets is prone to misconceptions common to conventional confidence intervals, such as misinterpreting the coverage rate and disregarding labels not included within the set.
6. Conclusion
Uncertainty quantification for deep Neural Networks (NNs) has been challenging due to their lack of interpretability. We study the conformal prediction set as an alternative solution to quantify the prediction uncertainty of NNs and whether communicating prediction uncertainty can help human decision-making in AI-advised image labeling. We contribute the results of a large pre-registered online experiment, in which we evaluate the utility of conformal prediction sets against status quo Top- predictions using stimuli that are varied by both in- and out-of-distribution (OOD), difficulty, and size. Our work shows that when the model is well-calibrated, a smaller prediction set is most beneficial, as larger sets can overwhelm users and decrease performance due to increased cognitive load. However, when the model faces unexpected OOD instances, large prediction sets can be more useful than their Top- counterparts. Our work sheds light on how conformal prediction sets can be used to rigorously quantify uncertainty for machine learning models in practical applications and outlines potential avenues for future research.
Appendix A Further Methodological Details
Corruption acc@1 acc@10 Coverage None 0.82 0.98 0.96 brightness 0.77 0.96 0.95 contrast 0.76 0.96 0.95 fog 0.72 0.93 0.94 jpeg compression 0.7 0.93 0.94 pixelate 0.66 0.9 0.92 motion blur 0.57 0.83 0.88 impulse noise 0.57 0.84 0.89 shot noise 0.56 0.83 0.89 gaussian noise 0.55 0.83 0.89 elastic transform 0.52 0.78 0.84 defocus blur 0.52 0.81 0.91 snow 0.51 0.79 0.86 frost 0.43 0.7 0.76 zoom blur 0.38 0.66 0.73 glass blur 0.29 0.55 0.72
Distribution of in-distribution cross-entropy loss with thresholds shown as vertical line
Corruption Size Difficulty Size1 Count SizeN Count Size Ratio SizeN Keep Size1 Include None small hard 539 2794 0.19 749 144 None small easy 13428 2962 4.53 1778 8060 defocus blur small hard 331 7750 0.04 870 37 defocus blur small easy 4571 2431 1.88 385 724 snow small hard 664 7319 0.09 827 75 snow small easy 4964 2554 1.94 451 877 frost small hard 993 9073 0.11 996 109 frost small easy 2896 1590 1.82 249 454 zoom blur small hard 722 9013 0.08 1066 85 zoom blur small easy 2925 1914 1.53 235 359 glass blur small hard 516 9973 0.05 1108 57 glass blur small easy 2025 1360 1.49 331 493
| Category | Open Codes | Description | ||
| 1 | UNINFORMATIVE | Content that doesn’t provide any significant or relevant information | ||
| 2 | HUMAN-AI COMPARISON | Participant analyzes differences and similarities between own’s decisions and AI predictions | ||
| 3 | HUMAN KNOWLEDGE | Participant makes decisions based solely on own intuition, wisdom, or experience | ||
| 4 | HUMAN KNOWLEDGE 1ST & AI ADVICE 2ND | Participant makes decision based on own’s intuition and experience first, supplemented by AI advice or suggestions | ||
| 5 | AI ADVICE 1ST & HUMAN CONFIRM 2ND | Participant seeks advice from AI first and then confirms with own’s knowledge | ||
| 6 | AI ADVICE WHEN HARD | Participant seeks advice from AI only and specifically in challenging or complex scenarios (e.g., blurry images, etc.) | ||
| 7 | TRUST AI PREDICTION(S) | Participant has confidence in the prediction made by AI | ||
| 8 | TRUST AI (RANKING OR CONFIDENCE) |
| 9 | TOP PREDICTION FALLACY | Participant expects to see the true answer in top predictions, but they believe the top prediction is wrong | ||
| 10 | DISAPPOINTED AT AI | Participant specifically mentions dissatisfaction with AI’s performance and does not trust AI’s prediction | ||
| 11 | SEARCH | Participant performs a general search without specifying | ||
| 12 | SEARCH (BY DROPDOWN) | Participant types keywords in the input field and refers to the dropdown for matches | ||
| 13 | SEARCH (BY KEYWORD) | Participant clicks “Search” to view a hierarchy tree that highlights matching categories | ||
| 14 | SEARCH (BY AI LABEL) | Participant clicks an AI-predicted label for a bottom-up search | ||
| 15 | SEARCH (HIERARCHY NETWORK) | Participant expands or subtracts nodes/categories and navigates through the provided hierarchy network | ||
| 16 | ELIMINATION | Participant tries to narrow down their options (to search in) by elimination | ||
| 17 | REPRESENTATIVE IMAGES | Participant compares the task image against the given label-representative images | ||
| 18 | COLOR & SURROUNDINGS | Participant relies on color, shapes, sizes, and surrounding objects to figure out the correct label |
Appendix B Additional Results
B.1. Shortest Path Length
The shortest path (SP) length quantifies the “incorrectness” of participants’ labeling choices relative to the correct label in the label space hierarchy network. Similar to accuracy, we analyze the variations of participants’ labeling errors by task image types. We report the median of expected predictions for each type of image stimuli with uncertainty quantified as 95% HDI, holding the trial effect constant at the mean.
B.1.1. In-distribution
At a high level, the SP length is, on average, smaller for easy images across conditions. When task images are hard, the SP length is greater for images with large set sizes across conditions.
When task images are easy, the SP length does not appear to be affected by set size—the distributions of expected SP length closely overlay across conditions (cf. Figure 7, (A) and (B)). After marginalizing by set sizes, we find that when participants have access to predictions, even when the provided label is wrong, the incorrect label appears to be closer to the correct label in distance (HPI: ) than those of the baseline condition (HPI: ).
When task images are hard but with a smaller average set size of , Figure 7 (C) shows the SP length is smallest for RAPS condition with a median of (HPI: ). We do not detect differences between the two Top- conditions, as HPIs closely overlap (Top-1: HPI: and Top-10: HPI: ). However, when without predictions, the baseline participants tend to submit more incorrect answers, reflected by the largest SP length of (HPI: ).
For hard task images with a larger average set size of , all intervals in Figure 7 (D) get wider than those in (C), indicating larger variations. Although SP length increased across conditions, we did not detect noticeable differences among conditions with predictions (HPI: ). Similarly, the baseline participants produce more inaccurate responses with a median SP length of (HPI: ).
In summary, we find that the SP length mirrors accuracy in that prediction display that yields higher accuracy tends to produce a lower SP length. Between accuracy and SP length, a noticeable difference can be detected from task images that are hard with a smaller set size: although participants in the Top-1 and the baseline conditions have similar labeling accuracy, the SP length of Top-1 participants is noticeably lower than that of the baseline. It implies that predictions can help elicit responses closer to the correct label, even if they are wrong. Otherwise, our SP results are mostly consistent with our accuracy results.
B.1.2. Out-of-distribution
For OOD task images, the SP length between easy and hard images becomes more contrasting. For easy images, because predictions are still reliable and accurate, we find that the Top-1 condition tends to produce the lowest SP length of (HPI: ) than Top-10 (; HPI: ) and RAPS (; HPI: ), after marginalizing over small and large set sizes (cf. Figure 7, (E) and (F)).
For hard OOD task images, we see an apparent mirroring effect relative to accuracy. When image stimuli have smaller average set size of shown in Figure 7 (G), Top-1 participants have the lowest accuracy rate and also provide labels that have the highest SP length of (HPI: ). RAPS participants have the highest accuracy for this type of image, produce the lowest SP length of (HPI: ). The Top-10 condition overlaps with RAPS with an SP length of (HPI: ), and both Top-10 and RAPS produce lower SP lengths than Top-1 and the baseline condition (; HPI: ).
Similarly, for hard OOD task images with a larger average set size of shown in Figure 7 (H), baseline and RAPS conditions that can help participants achieve the highest labeling accuracy, also support them to give answers closer to the correct label with a median SP length that ranges between and , followed by Top-1 (; HPI: ) and Top-10 conditions (; HPI: ).
However, despite that participants of the prediction set have similar accuracy for hard OOD images with small and large sets, we find SP length is smaller for images with larger set sizes. This implies that a larger set size is helpful for hard OOD images, guiding participants to the labels that are closer to the ground truth.
B.2. Distribution of Willingness-to-Pay
Condition Min. Q1 Q2 Q3 Max. Mean SD RAPS 0 1 1.0 2 8 1.78 1.74 Top-1 0 1 1.0 2 8 1.70 1.56 Top-10 0 1 1.5 2 8 1.84 1.67
Distribution of willingness-to-pay elicited from participants.
B.3. Qualitative Analysis of Strategies
Summarizing the qualitative results per identified open code.